# Decision-dependent Robust Charging Infrastructure Planning

### v1

@@ methods
Figure 15: Zone-wise differences in the number of slow and fast chargers in four stochastic cases compared to the benchmark
When comparing the three solution methods, the IPS method yields the highest number of chargers across all four instances using the full-year datasets. The DS and DS-h methods have an even lower number of chargers compared with the deterministic benchmark case, since an additional 18 trucks are excluded as HCVs due to the short parking duration. Fig. 14 shows the total number of chargers (i.e., the sum of fast and slow chargers) determined by the IPS procedure from monthly subproblems over 12 months in the four instances. From January to December, the number of trucks parking and charged in any of the eight zones rises from 156 to 182, leading to a steady increase in the required charger installations. As the parking duration uncertainty grows, the total number of EV chargers—particularly fast chargers—also increases. In the fourth instance, a larger share of fast chargers is installed compared to the third instance, since fast chargers are less affected by uncertainty than slow chargers. Across four stochastic instances, the only difference between the DS and DS-h results lies in the installation or placement of a single fast charger. This indicates that the assignment of specific fast chargers can hinder the convergence of the DS method. In contrast, the proposed heuristic prioritizes placing the fast charger, allowing the model to achieve the optimal solution more efficiently. The detailed solution process using the DS-h method is presented in Appendix 11.
Fig. 15 shows the zone-wise differences in the number of slow and fast chargers when comparing four stochastic instances to the benchmark case (i.e., case 1 in the deterministic formulation). In general, zones 1, 2, 4, and 7 have an increase in the number of slow chargers, while zone 6 has a reduction. It is because that zone 6 has a quite high variance in the parking duration. The model optimizes more chargers and schedules more trucks to be charged at zones 1 or 2, which are connected to zone 6, in four stochastic cases.
5.5.3 The impacts of uncertain parking duration
To evaluate the impact of uncertain parking duration on the charger planning and scheduling, we first simulated the monthly schedule based on the charger planning results for the case where $\sigma=2$ and $\gamma=0$, using the DS-h and IPS methods, respectively. We found that the charging installation obtained from the DS-h method meets the requirements for only six months, where the number of trucks dispatched is less, whereas those obtained from the IPS method satisfy the requirements for all 12 months. This discrepancy arises from the steadily increasing trend of truck parking and charging activities throughout the year, which the day-selection approach fails to capture adequately during peak months.
Instances
($\sigma$, $\gamma$)
Avg. slow
charging
hours
Avg. fast
charging
hours
Avg.
abandonment
hours
Avg. charging
hours
(truck per day)
Avg. charging
power (kW)
(0, 1) 3757.6 1280.7 20212.9 0.83 8.78
(1, 1) 2166.9 1249.9 20229.9 0.56 12.95
(2, 1) 2126.3 1357.5 20155.1 0.57 12.70
(2, 0.5) 515.1 1033.0 20191.6 0.25 28.59
Table 5: Summary of the average scheduled charging hours, abandonment hours, and charging power in a year across four instances
Figure 16: Boxplots of the monthly slow and fast charging hours across four stochastic instances versus the average of charging hours by month
We then evaluated four stochastic instances using the IPS method by simulating the charging scheduling in 12 months and calculated the average value of the total charging and abandonment hours by month. Fig. 16 shows the box-plots of the total number of slow and fast charging hours by months in four stochastic instances using the IPS method, and Table 5 summarizes the average scheduled charging hours, abandonment hours, and charging power by months. As the uncertainty increases where the standard deviation of the parking duration $\sigma$ increases from 0 to 2, the fast charger usage rises while the slow charger usage declines, since the shorter parking durations incentivize drivers to rely more on the fast charging. As the fast chargers are more frequently used, the average charging power increases from 8.78 kW when the standard deviation equals 0, to 12.95 kW and 12.70 kW when the standard deviation equals 1 and 2, respectively. The average abandonment hour keeps the same level across four stochastic instances.
For the second and third instances, where the standard deviations are 1 and 2, respectively, the third instance exhibits longer total charging hours and lower charging rates, even though fast-charging hours are greater than in the second instance. This is because the increased uncertainty in parking duration prolongs overall charging time, offsetting the benefit of more frequent fast charging. In contrast, in the fourth instance—where fast chargers effectively mitigate parking duration uncertainty—the total charging time decreases significantly, with trucks relying on fast chargers for most of the period. The average monthly slow-charging time is only 515.1 hours.
@@ 5.6 Conclusion and discussion
We developed a two-stage robust charging infrastructure planning model for the electrification of light-duty trucks at industrial sites with mixed-type chargers with different charging power. The first stage plans the charging infrastructure based on the selected charging locations and charging types. The second stage optimizes the charging schedules considering different waiting times, charging abandonment, and range anxiety. A case study is conducted on an open-pit mining site with eight potential charging zones. We used three input datasets extracted from the whole-year GPS data from around 200 trucks: (i) parking indications at each half-hour interval, (ii) actual parking durations of each parking window, and (iii) half-hour energy consumption of all trucks.
5.6.1 The impacts of overnight charging and range anxiety on charging behaviors
@@
Among eight zones, we assume that two designated zones allow overnight parking; otherwise, trucks will forgo charging if no chargers are available, unless their battery level falls below 30%. The battery SoC status below 30% is assumed to trigger the range anxiety, resulting in an additional half-hour waiting time. The benchmark case based on the selected 39 days yields 72 slow chargers and 6 fast chargers. We then examine three cases where overnight charging, range anxiety, and limited parking duration in each parking window are not considered, respectively. Extending the parking duration to the full window reduces the number of fast chargers to 3. Removing the designated overnight zones leads to a dramatic increase in abandonment hours by 1.63 times and an additional 20 slow chargers installed in total. No range anxiety requires seven more slow chargers to be installed. In addition, we found that the charging hours only occupy around 4% of the total parking time, meaning the limited waiting time and consequential abandonment are the main reasons that lead to a number of slow chargers at industrial sites.
5.6.2 Robust charging infrastructure planning considering uncertain parking duration
To utilize the full-year dataset, we proposed two methods—the DS and IPS approaches—for charging infrastructure planning. The uncertain parking duration within each half-hour parking window is represented using a robust uncertainty set, whose bounds depend on the charging decisions. The IPS method ensures robust scheduling performance throughout the year by planning for a larger number of chargers, whereas the DS method fails in several months when the charging demand peaks.
When the uncertainty of charging duration increases (i.e., the lower bound of the uncertainty set decreases), the total number of fast-charging hours rises, while the total number of slow-charging hours declines. Consequently, the overall average charging power increases under the mixed-type charging infrastructure. However, further reducing the lower bound of the robust set for parking duration to the minimum level (i.e., 5 minutes) leads to longer total charging hours and a decrease in average charging power. If the model incorporates the mitigating effect of fast charging on parking duration uncertainty—observed in many real-world studies—a smaller number of slow chargers would be planned and utilized.
5.6.3 Heuristic methods to assist the robust scheduling with a long optimization horizon
@@
The proposed two-stage planning and scheduling framework is formulated as a large-scale MILP model. For a 39-day scheduling instance with 189 trucks, the model contains over 1.5 million binary variables without pre-solving, making it computationally challenging. To address this, we introduce a fix-and-optimize heuristic for the charging scheduling under uncertainty. This is to forecast the high charging demand due to reduced parking duration and to pre-assign fast charging actions in the scheduling. This reduces the computation time by approximately 2-5 times for 39-day instances and yields the optimality gap of the problem below 0.1%.
@@
Finally, several promising avenues remain for future research. For example, future work could incorporate additional charging options, such as multiple types of fast chargers with rated powers exceeding 150 kW, to further reduce the number of required HCVs. Moreover, charging behavior modeling could be improved by integrating empirical data from existing electrified fleets to account for factors like variations in range anxiety at different times of day.

### v2

@@ 4.5.3 The impacts of different RO-DDU formulations for parking duration uncertainty
Figure 13: Zone-wise differences in the numbers of slow and fast chargers for three RO cases: RO - DI, RO-DDU (Instantaneous), and RO-DDU (Expect). The benchmark case is the RO - DI case
Sample Cases FC Hrs (truck $\cdot$ day) SC Hrs (truck $\cdot$ day) Ab Hrs (truck $\cdot$ day) Pct. of low SoC (%) Avg. CP (kW) Tot. Penalty (\$/truck $\cdot$ day)
In sample RO-DI 0.16 0.45 4.14 1.52 9.47 4.43
Instant. 0.22 0.52 4.13 1.72 7.80 4.51
Expect. 0.16 0.46 4.10 2.35 9.48 4.40
Out of sample Instant. 0.21 0.39 3.77 4.38 9.28 4.02
Expect. 0.21 0.39 3.77 4.90 9.33 4.01
Table 6: Summary of performance metrics for in-sample and out-of-sample evaluations under RO-DI, instantaneous DDU, and expectation-based DDU at the low-penalty case. Values are rounded to two decimal places. The values in bold indicate the highest value within each sample group and column.
Table 5 summarizes the charger installation results of three RO cases obtained in the in-sample datasets. Using the IPS approach can decompose the problem and solve three instances quickly, compared to the HS method. It relaxes the battery energy balance over a long scheduling horizon. Fig. 13 shows their zone-wise differences in the numbers of installed slow and fast chargers. Only when using the cluster formulation, the required number of chargers increases, particularly in zones 6, 7, and 8, where the charging duration variability is higher than in other zones. In addition, the cluster formulation requires substantially longer solution times due to the introduction of nonlinear charging-power constraints.
Table 6 summarizes the scheduling performance metrics of in-sample and out-of-sample tests in three cases. During in-sample tests, the RO-DI formulation yields the highest charging power, followed by RO-DDU (Expect) and RO-DDU (Instantaneous). This aligns with the theoretical analysis of charging power bounds via cluster-level aggregation in Section 7.3. The RO-DDU (expect) formulation accounts for decision dependence and the correlation among charging decisions within the same cluster. Therefore, the RO-DDU (expect) formulation has the highest scheduling penalty among the three cases. In out-of-sample tests, the scheduling penalty obtained from the RO-DDU (expect) formulation becomes less than the RO-DDU (Instantaneous) formulation. This means that, by accounting for correlations within the cluster, the expected formulation yields more efficient scheduling.
Data Range $\Delta_{c}$ Solving methods # of HCVs # of slow/fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
DS 0 HS 19 68/6 330,000 35,678 9.06 11.1
HS - h 19 68/6 330,000 35,679 0.07 2.40
DS $\pm\sigma$ HS 19 68/6 330,000 35,658 9.04 11.1
HS - h 19 69/6 331,500 36,065 0.05 2.15
DS $\pm\sigma$ HS 19 67/8 442,500 35,565 23.19 11.1
DDU HS - h 19 70/6 330,000 22,455 0.06 3.29
DS $\pm 2\sigma$ HS 19 68/6 330,000 35,736 9.02 11.1
HS - h 19 69/6 331,500 35,863 0.05 5.20
Table 7: Summary of optimization results with the varying size of uncertainty set based on the selected 39-day and full-year dataset using HS, HS with heuristic (HS - h)
Tackling high computation costs with the ‘fix-and-optimize’ heuristic: Considering the high computation costs of the 39-day instance under HS scheduling, we leverage the fix-and-optimize heuristic to solve problems. We impose a maximum time limit of 40,000 seconds for each instance and report the optimality gap when the solver fails to converge within this limit. In addition, we initialize the master problem as the warm start.
As presented in Table 7, we found that solving these HS instances directly is difficult, with all four cases reaching the maximum time limit. Using the proposed heuristic method (HS - h), the problem can reach an acceptable optimality gap of less than 0.1% while keeping the computation time within the prescribed limits. We used a warm start because the solution to the deterministic benchmark case is available, and completing 10 iterations takes approximately one hour. Across cases with varying uncertainty sets, the only difference between the HS and HS-h methods is the installation or placement of a single fast charger. This indicates that assigning specific fast chargers can hinder convergence. In contrast, the proposed heuristic prioritizes placing the fast charger, allowing the model to converge more efficiently. The detailed solution processes using the HS-h method are presented in Appendix 11.
4.5.4 Robust infrastructure planning based on the full year dataset
To ensure robust infrastructure planning over the year, we use the full-year dataset to plan the charging infrastructure under the IPS methods, as described in Section 3.1. These three full-year instances, corresponding to the size of uncertainty set increasing from 0 to $2\sigma$, provide the most robust planning results compared with the selective datasets.
Data Range $\Delta_{c}$ Solving methods # of HCVs # of slow/fast chargers Installation costs (\$) Scheduling penalty (\$) Gap (%) Time (Hrs)
FY 0 IPS 48 77/5 305,500 33,466 0.09 11.3
$\pm\sigma$ IPS 48 75/7 378,500 33,331 0.05 12.0
$\pm 2\sigma$ IPS 48 76/8 418,000 33,402 0.05 23.8
Table 8: Summary of optimization results with the full-year dataset using the IPS method. For the IPS method, the reported gap is the optimality gap of the final period in the iterative planning.
Figure 14: The total number of chargers in each month based on the full-year dataset using the IPS method in three instances with varying sizes of uncertainty sets (in colored lines) versus the total number of trucks parking and charged (in gray bars) in all eight zones
Table 7 summarizes the optimization results. When extending the data range to the full year using the IPS method, between 1 and 3 iterations exceeded the time limit. However, all three full-year instances converge by the end of the 12-month horizon. In the final instance involving joint planning and scheduling, an optimality gap of less than 0.1% is achieved. Fig.15 shows zone-wise differences in the number of slow and fast chargers in three full-year cases with the increasing size of uncertainty sets. As uncertainty increases, two or three additional fast chargers are replaced with slow chargers.
Figure 15: Zone-level variations in the number of slow and fast chargers across three full-year cases are illustrated. The benchmark corresponds to the case with a zero-sized uncertainty set, in which parking duration is fixed at the mean value for each cluster.
When comparing results from the 39-day dataset with those from the full-year dataset, the latter yields more chargers because the full-year data captures the end-of-year period with the highest vehicle counts. In contrast, the results derived from the 39-day dataset require even fewer chargers than the deterministic benchmark, because an additional 18 trucks are excluded from the HCV set due to their very short parking durations. Fig. 14 shows the total number of chargers (i.e., the sum of the number of fast and slow chargers) determined by the IPS procedure from monthly subproblems over 12 months in three instances. From January to December, the number of trucks parked and charged across the eight zones increases from 156 to 182, resulting in a steady rise in the number of required charger installations.
As the parking duration uncertainty grows, the total number of EV chargers increases, particularly for fast chargers. We then evaluated three full-year instances by calculating their performance metrics across 12 months. Table 9 summarizes the performance metrics of all three cases. As the uncertainty increases, both fast and low charger utilization rates reduce. The average abandonment hour remains constant across the three stochastic instances.
$\Delta_{c}$ FC Hrs (truck $\cdot$ day) FCUR (%) SC Hrs (truck $\cdot$ day) SCUR (%) Ab Hrs (truck $\cdot$ day) Pct. low SoC (%) Avg. CP (kW)
0 0.21 31.39 0.62 5.98 6.64 6.97 12.44
$\pm\sigma$ 0.21 21.88 0.36 3.54 6.64 2.46 13.18
$\pm 2\sigma$ 0.22 20.80 0.35 3.43 6.64 3.59 12.80
Table 9: Summary of all performance metrics in three stochastic cases. The number in bold highlights the highest values among the three cases.
Figure 16: (a) Distributions of monthly slow and fast charging hours of the entire truck fleet and (b) the monthly charging hour trends across a year in three uncertainty scenarios
Fig. 16 (a) shows distributions of monthly slow and fast charging hours of the entire truck fleet in the three stochastic instances, and Fig. 16 (b) shows the total monthly charging hour trends. By comparing three cases, we have the following observations: (1) With the increasing uncertainty of the charging duration, the total charging hours (the parking slots being decided to charge) reduce, but the fast charging hours increase. This indicates that the scheduling system seeks to avoid parking slots with highly uncertain parking durations and instead uses fast chargers more frequently. (2) If we further increase the uncertainty of charging duration, the total charging hours will increase, which is opposite to the previous trend, as more charging slots are required to meet the energy consumption.
@@ 4.6 Conclusion and discussion
We develop a robust model for planning and scheduling the charging infrastructure required to electrify light-duty trucks at industrial sites with mixed charger types. The model determines charging infrastructure deployment by selecting charging locations and charger types, and jointly optimizes charging schedules while accounting for waiting time, charging abandonment, and range anxiety. We conduct a case study at an open-pit mining site comprising eight charging zones. Three input datasets are constructed from year-long GPS records of approximately 200 operational trucks: (i) parking time slots discretized into half-hour intervals, (ii) realized parking durations within each half-hour window, and (iii) half-hourly energy consumption of all trucks. We assume fixed charging efficiencies for both slow and fast chargers and do not account for energy consumption during parking periods.
4.6.1 No overnight charging or no range anxiety will significantly increase the charger installations
@@
Among eight zones, we assume that two designated zones allow overnight parking; otherwise, trucks will forgo charging if no chargers are available. The battery SoC status below 30% is assumed to trigger the range anxiety, resulting in an additional half-hour waiting time. We found that neither overnight charging nor range anxiety significantly increases the number of charger installations, as both behaviors entail additional waiting time and reduce the need for installations. In addition, the charging hours account for only about 4% of the total parking time. This means limited waiting time and resulting abandonment are the main reasons for the number of charger installations at industrial sites.
4.6.2 A greater uncertainty in charging duration leads to changes in total charging hours as well as shifts in charging type
We assume that parking duration in the parking window is stochastic. To identify patterns, hierarchical clustering is used to group trucks. Then, the uncertain parking duration of each truck can be represented as a robust, feature-based uncertainty set. As charging duration uncertainty increases, total charging hours decline, reducing utilization rates for both slow and fast chargers. Increased uncertainty also shifts charging patterns, with total fast-charging hours increasing while slow-charging hours decrease. However, when the lower bound of the robust parking-duration set is further reduced to its minimum value (i.e., 5 minutes), the total charging hours increase to meet each truck’s energy requirements.
4.6.3 Alternative decision-dependent formulations that rely on clustering produce different charger planning results
We consider that the uncertainty of charging duration is decision-dependent, which impacts and is impacted by charging decisions (i.e., whether or not charging and charging types). We then investigate two decision-dependent formulations, assuming an affine dependence on instantaneous and cluster-level expected decisions. We found that cluster-level expected decisions yield tighter bounds on scheduling and a higher charging power level when we assume that charging decisions reduce parking duration. The instantaneous formulation always imposes a certain risk on individual parking time slots and neglects aggregated, correlated charging/parking behaviors across the fleet.
4.6.4 Fix-and-optimize heuristic can assist the robust scheduling depending on the forecasts
@@
The proposed planning and scheduling framework is formulated as a large-scale MILP model. For a 39-day instance with 188 trucks, the model contains over 1.5 million binary variables during pre-solving, making it computationally challenging. To address this, we introduce a fix-and-optimize heuristic for scheduling the charging operations under uncertainty. This is to forecast high charging demand due to reduced parking durations and to schedule fast-charging actions. This reduces the computation time by approximately 2-5 times for 39-day instances and yields the optimality gap of the problem below 0.1%.
@@
Finally, several promising avenues remain for future research. For example, future work could incorporate additional charging options, such as multiple types of fast chargers with rated powers exceeding 150 kW, to further reduce the number of required HCVs. Moreover, charging behavior modeling could be improved by integrating empirical data from existing electrified fleets to account for factors such as variations in range anxiety across times of day.